# Bridging Simplicity and Sophistication using GLinear: A Novel Architecture for Enhanced Time Series Prediction

### v2

@@ IV Methodology
  • It is a simpler model; it is not made up of any complex components, functions or blocks (like self-attention schemes, positional encoding blocks, etc.). It integrates two components: (1) a non-linear GeLU-based transformation layer to capture intricate patterns, and (2) Reversible Instance Normalization (RevIN).
@@
  • Due to its simple architecture, training of this model is very fast as compared to other transformer based predictors.
  • This proposed model provides comparable performance to other state-of-the-art predictors.
@@ V Experimental Setup
In this section, we present the details about used datasets, experimental setup, and evaluation metrics.
@@ V-A Dataset
We conducted experiments on four different real-world datasets (ETTh1, Electricity, Weather and Traffic). Table I provides a brief overview of these datasets.
@@ V-A1 ETTh1
The ETTh1 dataset is a multivariate time series dataset having 7 different variables (channels). It contains data of 725.83 days with granularity of 1 hour, meaning each timestamp represents a one-hour interval of data that provides 17420 timestamps values of each variable (17420 / 24 hours per day = 725.83 days of data).
@@ V-A2 Electricity
The Electricity dataset [7] is also a multivariate time series dataset having 321 channels. It contains data of 1096 days with granularity of 1 hour that provides 26304 timestamps values of each channel (26304 / 24 hours per day = 1096 days of data).
@@ V-A4 Traffic
The Traffic dataset [9] contains 731 days of data with a granularity of 1 hour. It provides data of 862 channels each having 17544 timestamps values.
@@ V-B Implementation Details
The GLinear model is implemented using Python and is sourced from the official Pytorch implementation of LTSF-Linear  [^1] https://github.com/cure-lab/LTSF-Linear/. Respective code repository contains training and evaluation protocol of Autoformer, NLinear and DLinear. Similarly, the RLinear  [^2] https://github.com/plumprc/RTSF/blob/main/models/RLinear.py model is trained and evaluated using the same protocol to ensure a fair comparison across all models.
The same set of hyperparameters are used for training all linear models, such as using the Mean Squared Error (MSE) criterion , Adam [11] optimizer and a learning rate of 0.001.
@@ VI Experiments and Results
TABLE II: Multivariate Forecasting performance comparison across datasets and models (lower MSE/MAE is better). Input Sequence Length for all experiments is 336 and a range of Prediction Length {12, 24, 48, 96, 192, 336, 720} are used. The top-performing results are marked in bold. The second-best results are underlined.
Lookup Window (Input Sequence Length) = 336 Learning Rate 0.001
@@
Dataset / Output Horizon (Prediction Length) MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
@@
Figure 5: Multivariate Forecasting performance comparison across datasets and models. Input Sequence Length is 336 and values of Prediction Length are {12, 24, 48, 96, 192, 336, 720}. Datasets are (a) Electricity, (b) ETTh1, (c) Traffic, and (d) Weather.
@@ VI-A Evaluating Different Prediction Lengths for Fixed Input Length
In the first experiment, the length of the input sequence was fixed at 336 time steps, representing a historical window to learn the underlying patterns. The prediction lengths were varied across multiple time frames to assess the model’s ability to forecast different horizons, as displayed in Table II. Table II provides a comprehensive evaluation of all predictors, including Autoformer, NLinear, DLinear, RLinear, and the proposed GLinear across four datasets. In an extensive series of experiments, we varied the prediction lengths in range: {12, 24, 48, 96, 192, 336, 720}. This set of experiments allows for an evaluation of the model performance in forecasting short, medium and long-term future steps, helping to gauge the effectiveness of the model for various forecasting time horizons. We used a quantitative approach to score the best-performing candidate in all datasets, giving a point for the best and second-best results in each row. The results highlight that Glinear is a majority winner with substentially improved results in both evaluation metrics.
The performance of these models can also be compared using Figure 5. The results demonstrate that the proposed GLinear model outperforms other predictors in most cases (a lower MSE indicates better performance). Specifically, GLinear achieves the highest performance for the Electricity and Traffic datasets. For Weather forecasting, GLinear is the second-best model, with RLinear taking the top spot. However, the difference in MSE between the two models is minimal, as shown in Figure 5 (d). For the ETTh1 dataset, no single model consistently outperforms others across all prediction lengths. For the ETTh1 dataset, GLinear ranks among the top-performing models for shorter prediction lengths (12, 24, and 48), but the best-performing model changes as the prediction length increases. Additionally, the performance of GLinear deteriorates with longer prediction lengths for this dataset. The next experiment aims to provide an explanation to better understand the underlying reasons.
@@ VI-B Impact of Input Sequence Length on Forecasting Future Steps
Figure 6: A performance comparison of all linear predictors (NLinear, DLinear, RLinear, and Proposed GLinear) with varying input sequence lengths in both long-term forecasting (720 time steps) and short-term forecasting (24 time steps) for different datasets, (a) Electricity, (b) ETTh1, (c) Traffic, and (d) Weather.
@@
Figure 6 provides the result of the second experiment, which compares the performance of proposed GLinear with other state-of-the-art Linear predictors (NLinear, DLinear and RLinear) under different input prediction lengths to understand how much historical data is enough for short- and long-term forecasting. MSE results were computed for two prediction lengths (24 and 720 time steps) to analyze short-term and long-term forecasting performance, respectively.
@@
The performance comparison of predictors on the ETTh1 dataset reveals interesting insights about the proposed GLinear model. Specifically, GLinear’s performance declines as it uses more historical data. However, it delivers the best performance for both short- and long-term forecasting when using shorter input sequence lengths, making it data-efficient by requiring less historical data to produce accurate forecasts.
@@ VII Conclusion and Future Directions
In this paper, we have introduced GLinear, a novel and data-efficient architecture for multivariate time series forecasting (TSF) that leverages periodic patterns to enhance prediction accuracy while requiring less historical data compared to existing linear models. Our experiments across four datasets—ETTh1, Electricity, Traffic, and Weather—demonstrate that GLinear not only achieves competitive performance but also outperforms state-of-the-art models such as NLinear, DLinear, and RLinear, as well as Transformer-based models like Autoformer, in various TSF scenarios.
@@
Future work can explore applying the GLinear architecture to other time series-related tasks, such as anomaly detection or forecasting in different domains. Additionally, the periodic pattern extraction mechanism can be integrated into other deep learning models to enhance their efficiency and predictive performance.
Acknowledgment
This research has been partially funded by the European Union - Next Generation EU through the Project of National Relevance ”Innovative mathematical modeling for cell mechanics: global approach from micro-scale models to experimental validation integrated by reinforcement learning”, financed by European Union-Next-GenerationEU-National Recovery and Resilience Plan-NRRP-M4C1-I 1.1, CALL PRIN 2022 PNRR D.D. 1409 14-09-2022—(Project code P2022MXCJ2, CUP F53D23010080001) granted by the Italian MUR.

### v4

@@ IV-B Features of GLinear
  • It is a simpler model; it is not made up of any complex components, functions, or blocks (like self-attention schemes, positional encoding blocks, etc.). It integrates two components: (1) a non-linear GELU-based transformation layer to capture intricate patterns and (2) Reversible Instance Normalization (RevIN).
@@
  • Due to its simple architecture, training of this model is significantly faster than transformer-based predictors. The parameter count per channel is
    $|W^{(1)}|+|b^{(1)}|+|W^{(2)}|+|b^{(2)}|=L^{2}+L+TL+T.$ (8)
    With $C$ channels (independent weights), this scales as $O(C\cdot(L^{2}+TL))$. This is significantly more efficient than transformer-based models, which scale as $O(L^{2}C)$ in the self-attention layers.
    GLinear has the potential to provide competitive performance to other state-of-the-art linear predictors over different benchmarks. Unlike NLinear, which only tackles distribution shift, and DLinear, which requires explicit trend decomposition, GLinear introduces smooth Gaussian nonlinearity within a lightweight linear framework. This enables it to capture nonlinear seasonality while retaining the computational efficiency of linear predictors, bridging the gap between purely linear mappings and heavy transformer-based models.
@@ V Experimental Setup
In this section, we present the details about the datasets used, the experimental setup, and the evaluation metrics.
@@ V-A Dataset
We conducted experiments on four different real-world datasets (ETTh1, Electricity, Weather, and Traffic). Table I provides a brief overview of these datasets.
@@ V-A1 ETTh1
The ETTh1 dataset is a multivariate time series dataset having 7 different variables (channels). It contains data of 725.83 days with a granularity of 1 hour, meaning each timestamp represents a one-hour interval of data that provides 17420 timestamp values of each variable (17420 / 24 hours per day = 725.83 days of data).
@@ V-A2 Electricity
The Electricity dataset [35] is also a multivariate time series dataset having 321 channels. It contains data of 1096 days with a granularity of 1 hour, which provides 26304 timestamp values of each channel (26304 / 24 hours per day = 1096 days of data).
@@ V-A4 Traffic
The Traffic dataset [37] contains 731 days of data with a granularity of 1 hour. It provides data from 862 channels, each having 17,544 timestamp values.
@@ V-B Implementation Details
The GLinear model is implemented using Python and is sourced from the official PyTorch implementation of LTSF-Linear  [^2] https://github.com/cure-lab/LTSF-Linear/. The respective code repositories contain the training and evaluation protocols of Autoformer, NLinear, and DLinear. Similarly, the RLinear  [^3] https://github.com/plumprc/RTSF/blob/main/models/RLinear.py model is trained and evaluated using the same protocol to ensure a fair comparison across all models.
The same set of hyperparameters is used for training all linear models, such as using the Mean Squared Error (MSE) criterion, the Adam [38] optimizer, and a learning rate of 0.001. The source code is publicly available on GitHub.
@@ VI Experiments and Results
TABLE II: Multivariate forecasting performance comparison across the datasets and models (lower MSE/MAE is better). The input sequence length for all experiments is 336 time steps, and a range of prediction lengths {12, 24, 48, 96, 192, 336, 720} is used. The top-performing results are marked in bold. The second-best results are underlined.
Lookup Window - Learning Rate 0.001 (Input Sequence Length) = 336
@@
Dataset/Output Hori. (Prediction Length) MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
@@
Figure 5: Multivariate Forecasting performance comparison across the datasets and the models. Input sequence length is 336, and values of prediction lengths are {12, 24, 48, 96, 192, 336, 720}. The datasets are (a) Electricity, (b) ETTh1, (c) Traffic, and (d) Weather.
TABLE III: The average per-epoch training time and estimated per-sample inference latency on the Electricity dataset (Input length = 336, prediction length = 720, batch size = 16). Inference latency is estimated from average iteration speed in the training logs.
Model Per-epoch Training Time (s) Inference (ms/sample)
NLinear [12] 21.32 2.01
DLinear [12] 24.75 2.27
RLinear [14] 233.10 15.83
GLinear (ours) 23.22 2.15
@@
For a fair comparison, we have focused on comparing GLinear against other linear models from the literature because these previous works (NLinear [12], DLinear [12], RLinear [14]) have already been benchmarked against five recent transformer-based methods (Autoformer [17], FEDformer [22], Informer [34], Pyraformer [39], and LogTrans [40]) and demonstrated that linear models outperform these transformers. To provide context while keeping the experiments concise, we compared GLinear with the linear baselines and Autoformer as a representative transformer.
@@ VI-A Evaluating Different Prediction Lengths for Fixed Input Length
In the first experiment, the length of the input sequence was fixed at 336 time steps, representing a historical window to learn the underlying patterns. The prediction lengths were varied across multiple time frames to assess the model’s ability to forecast different horizons, as displayed in Table II. Table II provides a comprehensive evaluation of all predictors, including NLinear [12], DLinear [12], RLinear [14], Autoformer [17], and the proposed GLinear across four datasets. In an extensive series of experiments, we varied the prediction lengths in the range: {12, 24, 48, 96, 192, 336, 720}. This set of experiments allows for an evaluation of the model’s performance in forecasting short-, medium-, and long-term future steps, helping to gauge the effectiveness of the model for various forecasting time horizons. We used a quantitative approach to score the best-performing candidate in all datasets, giving a point for the best and second-best results in each row. The results conclude that GLinear is a majority winner with substantially improved results in both evaluation metrics.
Compared strictly against the linear models (NLinear [12], DLinear [12], and RLinear [14]), smooth nonlinearities introduced by the GLinear architecture capture more complex patterns such as seasonality and nonlinear dependencies. By restricting operations to per-channel projections, GLinear preserves temporal locality and avoids spurious inter-channel correlations that often hamper the generalization of other linear models. The insertion of a light nonlinearity (GELU) endows the model with the capacity to approximate smooth nonlinear mappings, which pure linear predictors lack. Together with RevIN, which stabilizes scale shifts across instances, this design biases GLinear toward robust short-to-medium horizon forecasting. These inductive biases explain why GLinear can outperform both heavier Transformer-based models and simpler linear models across diverse datasets. In addition, GLinear’s per‑channel linear mapping with a single smooth nonlinearity, together with RevIN, surfaces component effects and instance‑level scale shifts more directly than deeper attention stacks (as deployed in the timeseries Transformers), which supports interpretability in multivariate settings. By avoiding quadratic attention and large attention blocks typical of time‑series Transformers, GLinear scales favorably in both training and inference; this aligns with Table III and prior works [12, 14] advocating lightweight mappings for long sequences.
Figure 5 further compares the forecasting performance of these models. The results demonstrate that the proposed GLinear model outperforms other predictors in most cases (a lower MSE indicates better performance). Specifically, GLinear achieves the highest performance for the Electricity and Traffic datasets. For weather forecasting, GLinear is the second-best model, with RLinear taking the top spot. However, the difference in MSE between the two models is minimal, as shown in Figure 5 (d). For the ETTh1 dataset, no single model consistently outperforms others across all prediction lengths. For the ETTh1 dataset, GLinear ranks among the top-performing models for shorter prediction lengths (12, 24, and 48), but the best-performing model changes as the prediction length increases. We noted that on ETTh1, GLinear’s performance degrades as the horizon lengthens. This effect is expected, since the channel-wise projection does not explicitly model long-range temporal dependencies, making the approximation error accumulate over larger horizons.
[…]
@@ VI-B Impact of Input Sequence Length on Forecasting Future Steps
Figure 6: A performance comparison of the linear predictors (NLinear, DLinear, RLinear, and the proposed GLinear) with varying input sequence lengths in both long-term forecasting (720 time steps) and short-term forecasting (24 time steps) for different datasets: (a) Electricity, (b) ETTh1, (c) Traffic, and (d) Weather.
@@
Figure 6 provides the result of the second experiment, which compares the performance of the proposed GLinear with other state-of-the-art linear predictors (NLinear [12], DLinear [12], and RLinear [14]) under different input prediction lengths to understand how much historical data is enough for short- and long-term forecasting. MSE results were computed for two prediction lengths (24 and 720 time steps) to analyze short-term and long-term forecasting performance, respectively.
@@
The performance comparison of predictors on the ETTh1 dataset reveals interesting insights about the proposed GLinear model. Specifically, GLinear’s performance declines as it uses more historical data. This behavior is expected, as GLinear processes each channel independently with a compact two-layer MLP, which excels at modeling local temporal patterns but has limited capacity to capture very long-term dependencies. Across the four datasets and horizons evaluated, GLinear achieves competitive performance and often outperforms NLinear, DLinear, RLinear, and the representative Transformer baseline (Autoformer), while exhibiting efficient training and inference.
@@ VII Conclusion and Future Directions
In this paper, we have introduced GLinear, a novel and data-efficient architecture for multivariate time series forecasting (TSF) that leverages periodic patterns to enhance prediction accuracy while requiring less historical data compared to existing linear models. Our experiments across four datasets (ETTh1, Electricity, Traffic, and Weather) demonstrate that GLinear not only achieves competitive performance but also outperforms state-of-the-art models such as NLinear, DLinear, and RLinear, as well as Transformer-based models like Autoformer, in various TSF scenarios.
@@
Future work could explore enhancements such as dilated convolutions, lightweight attention modules, or temporal decomposition to improve long-term forecasting performance, and apply the GLinear architecture to other time series-related tasks, such as anomaly detection or forecasting in different domains. Additionally, the periodic pattern extraction mechanism can be integrated into other deep learning models to enhance their efficiency and predictive performance.